### v1

# Trio-ViT: Post-Training Quantization and Acceleration for Softmax-Free Efficient Vision Transformer

Abstract
Motivated by the huge success of Transformers in the field of natural language processing (NLP), Vision Transformers (ViTs) have been rapidly developed and achieved remarkable performance in various computer vision tasks. However, their huge model sizes and intensive computations hinder ViTs’ deployment on embedded devices, calling for effective model compression methods, such as quantization. Unfortunately, due to the existence of hardware-unfriendly and quantization-sensitive non-linear operations, particularly Softmax, it is non-trivial to completely quantize all operations in ViTs, yielding either significant accuracy drops or non-negligible hardware costs. In response to challenges associated with standard ViTs, we focus our attention towards the quantization and acceleration for efficient ViTs, which not only eliminate the troublesome Softmax but also integrate linear attention with low computational complexity, and propose Trio-ViT accordingly. Specifically, at the algorithm level, we develop a tailored post-training quantization engine taking the unique activation distributions of Softmax-free efficient ViTs into full consideration, aiming to boost quantization accuracy. Furthermore, at the hardware level, we build an accelerator dedicated to the specific Convolution-Transformer hybrid architecture of efficient ViTs, thereby enhancing hardware efficiency. Extensive experimental results consistently prove the effectiveness of our Trio-ViT framework. Particularly, we can gain up to and FPS under comparable accuracy over state-of-the-art ViT accelerators, as well as and DSP efficiency. Codes will be released publicly upon acceptance.
Index Terms:
Post-training quantization, hardware acceleration, Softmax-free efficient Vision Transformer.I Introduction
Thanks to the powerful global information extraction capability of self-attention mechanism, Transformers have achieved great success in various natural language processing (NLP) tasks [1, 2, 3]. This success has inspired the rapid development of Vision Transformers (ViTs) [4, 5], which have gained increasing attention in the field of computer vision and shown superior performance compared to their convolution-based counterparts. However, their enormous model sizes and intensive computations challenge the deployment of ViTs on embedded/mobile devices, where both memory and computing resources are limited. For example, ViT-Large [4] contains M parameters and yields G FLOPs during inference. Thus, effective model compression techniques are highly desired to facilitate ViT’s real-world applications.
Among them, model quantization [6, 7, 8] stands out as one of the most effective and widely adopted compression methods. It converts floating-point weights/activations into integers, leading to a reduction in both memory consumption and computational costs during inference. Unfortunately, due to the existence of non-linear operations, including LayerNorm (LN), GELU, and especially Softmax, which are not only hardware unfriendly but also quantization-sensitive, ViTs are difficult to be fully quantized, yielding either significant accuracy drops or notable hardware overhead [9, 10]. To solve these challenges, several efforts [9, 11, 10] have been devoted. For example, FQ-ViT [9] identifies extreme inter-channel variations in LN’s inputs and excessive non-uniform distributions in attention maps, and proposes Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS) for LN and Softmax quantization, respectively. Furthermore, I-ViT [11] develops innovative lightweight dyadic arithmetic methods to approximate ViTs’ non-linear operations, thus achieving integer-only inference. Despite their effectiveness, they are dedicated to the quantization for standard ViTs while overlooking the inherent quantization and acceleration opportunities within efficient ViTs [12, 13, 14], where the vanilla Softmax-based attention with quadratic computational complexity is typically traded with more efficient Softmax-free attentions that exhibit linear computational complexity. Thus, we redirect our focus towards the exploration of effective quantization and acceleration for efficient ViTs, aiming to fully unleash their potential algorithmic benefits to win both accuracy and hardware efficiency, i.e., (i) the Softmax-free property to boost the achievable quantization performance and (ii) the linear complexity characteristic of attention to enhance inference efficiency.
In addition to the algorithm level, various works have built dedicated accelerators to boost ViTs’ hardware efficiency from the hardware perspective [15, 16, 17]. For instance, Auto-ViT-Acc [16] adopts mixed quantization schemes, e.g., fixed-point and power-of-two, to quantize ViTs, and develops a dedicated accelerator to fully leverage the computational resources available on FPGAs. Moreover, ViTCoD [17] proposes pruning and polarization techniques to transform ViTs’ attention maps into denser and sparser variants, and then develops a dedicated accelerator incorporating both dense and sparse engines to simultaneously execute the above two workloads. Despite the superiority of the above accelerators in enhancing hardware efficiency, they are dedicated to standard ViTs and fall short in fully accelerating efficient ViTs [12, 13, 14], which are typically characterized by (i) Softmax-free linear attentions and (ii) Convolution-Transformer hybrid architectures. Specifically, it has been widely verified that the computational complexity reduction of linear attention yields a degradation in its local feature extraction ability, thus necessitating extra compensation components such as convolutions [12, 13]. This results in hybrid architectures for efficient ViTs that comprise both convolutions and Transformer blocks, thus calling for dedicated accelerators to unleash their potential benefits.
To grasp the inherent quantization and acceleration opportunities in efficient ViTs, we make the following contributions:
We propose Trio-ViT, a post-training quantization and acceleration framework for efficient Vision Transformers (ViTs) via algorithm and hardware co-design. To the best of our knowledge, this is the first work dedicated to the quantization and acceleration of efficient ViTs.
At the algorithm level, we conduct a comprehensive analysis of distinct activations of Softmax-free efficient ViTs and unveil specific quantization challenges. Then, we develop a tailored post-training quantization engine that incorporates several novel strategies, including channel-wise migration, filter-wise shifting, and log2 quantization, to address the involved challenges with boosted quantization accuracy.
At the hardware level, we advocate a hybrid design incorporating multiple computing cores to effectively support various operation types in the Convolution-Transformer hybrid architecture of efficient ViTs. Besides, we propose a pipeline architecture to facilitate both inter- and intra-layer fusions, thus enhancing hardware utilization and easing the bandwidth requirement.
Extensive experiments and ablation studies consistently validate the effectiveness of our Trio-ViT framework. For example, we can offer up to and FPS with comparable accuracy over state-of-the-art (SOTA) ViT accelerators. Besides, we can achieve up to and DSP efficiency. It is expected that our work can open up an exciting perspective for the quantization and acceleration of Softmax-free efficient ViTs.
The rest of this paper is organized as follows: we first introduce related works in Sec. II and preliminaries in Sec. III; Then Sec. IV and Sec. V separately illustrates Trio-ViT’s post-training quantization and dedicated accelerator; Furthermore, extensive experiments and ablation studies consistently demonstrates our Trio-ViT’s effectiveness in Sec. VI; Finally, Sec. VII summarizes this paper.
II Related Works
II-A Model Quantization for Vision Transformers (ViTs)
Model quantization, which represents floating-point weights and activations with integers without modifying model architectures, is a generic compression solution. It can be roughly categorized into two approaches: quantization-aware training (QAT) and post-training quantization (PTQ). Specifically, QAT [18, 8, 19, 11] involves weight fine-tuning to facilitate quantization, yielding higher accuracy or lower quantization bit. In contrast, PTQ [9, 10, 20, 21], which eliminates resource-intensive fine-tuning and streamlines models’ deployment, has recently gained increasing attention. For example, [20] incorporates an innovative ranking loss to preserve the functionality of the self-attention mechanism during quantization, successfully quantizing linear operations (matrix multiplications) in ViTs. Additionally, FQ-ViT [9] further introduces Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS) to quantize the hardware- and quantization-unfriendly non-linear operations (i.e., LayerNorm and Softmax) in ViTs, achieving full quantization. However, these works are developed for standard ViTs and cannot capture quantization opportunities offered by efficient ViTs [12, 13], which feature Softmax-free attention with linear computational complexity to win both quantization accuracy and hardware efficiency.
II-B Efficient ViTs
ViTs [4, 5, 22, 12, 13, 23, 24] have gained growing attention recently and have been developed rapidly in the computer vision field. Among them, ViT [4] firstly applies a pure Transformer to process sequences of image patches, achieving remarkable performance. Furthermore, DeiT [5] offers a better training recipe for ViT, significantly reducing training costs. However, ViTs still incur intensive computational costs and huge memory footprints during inference to achieve superior performance, calling for efficient ViTs [22, 12, 13, 23, 24]. Particularly, EfficientViT [12], the SOTA one, replaces the vanilla self-attention of quadratic complexity with a novel lightweight multi-scale attention, achieving a global receptive field while enhancing hardware efficiency. Besides, Flatten Transformer [13] opts for an innovative focused linear attention, preserving expressiveness with low computational complexity. Despite the inherent algorithmic benefits of Softmax-free linear attentions in efficient ViTs [12, 13], including (i) the Softmax-free property to facilitate quantization and (ii) the linear complexity to boost hardware efficiency, their dedicated quantization and acceleration methods remain under-explored.
II-C Transformer Accelerators
Recently, various works [25, 26, 16, 27, 17, 15] have developed dedicated accelerators to promote Transformers’ real-world deployment. Specifically, Sanger [25] dynamically prunes attention maps during inference and builds a reconfigurable accelerator with a score-stationary dataflow to accelerate such sparse patterns. As for ViT accelerators, VAQF [15] is designed to accelerate ViTs with binary weights and mixed-precision activations. Auto-ViT-Acc [16] incorporates heterogeneous computing resources available on FPGAs (i.e., DSPs and LUTs) to separately accelerate the mixed quantization scheme (i.e., fixed-point and power-of-two) for ViTs. ViTCoD [17] prunes and polarizes ViTs’ attention maps into denser and sparser ones, and constructs an accelerator to execute them on separate computing engines. While these methods can enhance the hardware efficiency for standard ViTs, they are not directly applicable for efficient ViTs [12, 13] due to the latter’s distinct model architectures, such as Softmax-free linear attention and Convolution-Transformer hybrid structures, calling for dedicated accelerators.
III Preliminaries
III-A Structure of Standard ViTs
As depicted in Fig. 1, input images are initially partitioned into fixed-size patches and further enhanced with token and positional embeddings, serving as input tokens for ViTs’ Transformer blocks. Each Transformer block comprises two key components: a Multi-Head Self-Attention module (MHSA) and a Multi-Layer Perceptron (MLP), both preceded by LayerNorm (LN) and followed by residual connections. Specifically, the MHSA is the core element in Transformers for global information capture. It projects input tokens into queries , keys , and values following Eq. (1), where , , and are corresponding weights for the head. Subsequently, as formulated in Eq. (2), is multiplied by the transposed () and then subjected to Softmax normalization (where represents the feature dimension of each head) to generate the attention map. This attention map is further multiplied by to obtain the attention output for the head. Finally, the attention outputs from all heads are concatenated and projected using weights to produce the final MHSA output, i.e., in Eq. (3). As for the MLP, it comprises two linear layers separated by the GELU activation function.
Limitations. There are two limitations of standard ViTs: (i) the quadratic computational complexity of the self-attention w.r.t. token numbers [22, 12, 13] and (ii) the hardware- and quantization-unfriendly non-linear operations, i.e., LN, GELU, and especially Softmax [9, 10], hindering ViTs’ achievable hardware efficiency and quantization accuracy.
III-B Structure of EfficientViT
To tackle the above limitations, efficient ViTs [12, 13, 27] have emerged as a promising solution. Here we take EfficientViT [12], the SOTA efficient ViT, as the example for illustration. As shown in Fig. 2, it not only (i) incorporates Softmax-free linear attention but also (ii) replaces vanilla LN and GELU with hardware- and quantization-friendly BatchNorm (BN) and Hardswish (Hswish) [29], respectively, significantly facilitating quantization and acceleration.
Specifically, EfficientViT [12] mainly consists of two types of blocks: MBConvs [28] and EfficientViT modules. Each MBConv comprises two point-wise convolutions (PWConvs) divided by a depthwise convolution (DWConv), each followed by a BN and a Hswish (except the last layer). Particularly, BN can be implemented using convolutions and seamlessly folded into preceding convolutions, simplifying quantization and acceleration [30]. Additionally, each EfficientViT module includes a lightweight Multi-Scale Attention (MSA) for context information extraction and an MBConv for local information extraction. In MSA, inputs are projected to generate , which are then processed by lightweight small-kernel convolutions to generate multi-scale tokens. After applying ReLU-based global attention, the results are concatenated and projected to produce final outputs. Notably, ReLU-based global attention essentially replaces the similarity function in vanilla Softmax-based attention with , thus allowing for (i) the removal of Softmax and (ii) the utilization of the associative property of matrix multiplication to reduce computational complexity from quadratic to linear. This reformulates Eq. (2) as follows:
IV Trio-ViT’s Post-Training Quantization
As illustrated above, due to the inherent benefits of the SOTA efficient ViT, dubbed EffcientViT [12], i.e., the vanilla Softmax-based self-attention with quadratic complexity, LN, and GELU, are replaced with hardware- and quantization-friendly ReLU-based global attention with linear complexity, BN, and Hardswish, respectively, we thus explore quantization and acceleration on top of EfficientViT to win both quantization accuracy and hardware efficiency.
We adopt the most widely applied hardware-friendly quantization setting [6] by default, i.e., the symmetric layer-wise and filter-wise uniform quantization for activations and weights , respectively. Formally, as expressed in Eq. (5), / are quantized /, / are the corresponding scaling factors, means rounding to the nearest, and is the quantization bandwidth. Particularly, we follow the SOTA PTQ method BRECQ [6], which uses the diagonal Fisher Information Matrix (FIM) to sequentially reconstruct basic blocks (e.g., MBConvs and Lightweight MSAs in EfficientViT), thus enhancing cross-layer dependency while maintaining generalizability. Given the FIM as the objective function, quantization of weights and activations are optimized via Adaround [31] and Learned Step size Quantization (LSQ) [32], respectively.
IV-A Observations
As we adopt the block-wise reconstruction following [6] for quantization optimization and MBConvs/lightweight MSAs are two primary blocks in EfficientViT, we begin by retaining matrix-multiplications (MatMuls) within MSAs in Eq. (4) at full precision to assess quantization impact on MBConvs.
IV-A1 Observations on Quantization of MBConvs
Although it has been widely recognized that activations are more sensitive to quantization than weights [9, 21], this sensitivity is exacerbated in the context of EfficientViT. As presented in Table I, quantizing only weights in EfficientViT-B1 [12] to -bit (W8) leads to a comparable accuracy () compared to its full-precision counterpart, while quantizing both weights and activations (except MatMuls in MSAs) to the same bit (W8A8) yields a catastrophic accuracy drop of . This emphasizes the extreme quantization sensitivity of activations in EfficientViT, particularly those in MBConvs, as evident in the accuracy comparison between columns and in Table II. As previously introduced in Sec. III-B, each MBConv contains two PWConvs separated by a DWConv, we then conduct an ablation study to individually quantize input activations of these three layers in all MBConvs. As shown in Table II, the input activations of DWConvs (DW) and the second PWConvs (PW2) are the most quantization sensitive and should be primarily responsible for the accuracy drop. To comprehend this issue, we visualize their input activations in Fig. 3 and observe two challenges.
| EfficientViT-B1[12] | W8 | W8A16 | W8A12 | W8A10 | W8A8 |
| Top-1 Accuracy* (%) | 79.39 | 79.32 | 78.86 | 75.08 | 3.23 |
| Drop (%) | 0.01 | 0.06 | 0.52 | 4.30 | 76.15 |
Tested on ImageNet with the input size of by default.
| EfficientViT-B1 | Head*+MBConvs | Head* | Head*+PW1 | Head*+DW | Head*+PW2 |
| Accuracy (%) | 3.23 | 79.24 | 79.13 | 28.72 | 8.85 |
| Drop (%) | 76.15 | 0.15 | 0.26 | 50.68 | 70.550 |
“Stem+Head” is abbreviated as “Head” here for simplicity.
Challenge # 1: Inter-Channel Variations in DW’s Inputs. Specifically, as shown in Fig. 3 (c), the input activation of DW exhibits significant inter-channel variations, resulting in the majority of values being represented with few quantization bins (see Fig. 3 (d)). For example, in the input activation of DW within the MBConv from the last stage in EfficientViT-B1, approximately values occupy a mere of total quantization bins. In contrast, this percentage is considerably higher at for PW1’s input, which is greater.
Challenge # 2: Inter-Channel Asymmetries in PW2’s Inputs. As depicted in Fig. 3 (e), input activations of PW2 feature extreme inter-channel asymmetries compared to those of PW1 in Fig. 3 (a), yielding a broader value range and thus a lower quantization resolution. For instance, in PW2’s input within the MBConv from the last stage of EfficientViT-B1, the interval of the first channel is (, ), while the interval among all channels is (, ), which is larger.
IV-A2 Observations on Quantization of Lightweight MSAs
When quantizing MatMuls in Eq. (4) within lightweight MSAs to 8-bit, we encounter notably worse results, which manifest as a “Not-a-Number” (NaN) issue. We find that this issue primarily arises from the quantization of divisors/denominators. As depicted in Figs. 4 (a) and (b), the wide range of values within divisors results in a reduced quantization resolution for smaller values when adopting the uniform quantization. Nevertheless, smaller values within divisors exhibit greater sensitivity compared to larger values. For instance, rounding a divisor of to during quantization results in an absolute difference of , yet it only doubles the final division results. In contrast, approximating the divisor of to incurs a mere absolute difference but increases the final results by a factor of . Particularly, changing a divisor of to yields a negligible absolute difference, but it causes the final results larger. These examples distinctly underscore the inherent incompatibility of uniform quantization for divisors, especially those exhibiting a wide range of values.
IV-B Channel-Wise Migration for DW’s Inputs
As introduced in Challenge # 2 in Sec. IV-A1, there exist extreme inter-channel variations in DW’s inputs, making the vanilla layer-wise quantization unsuitable. Fortunately, owing to the distinctive algorithmic characteristic of DWConvs, which process each input channel independently and thus eliminates the summations along channels, we can directly employ channel-wise quantization to assign individual scaling factors for each input channel. This approach effectively addresses the above challenge without compromising hardware efficiency. Nonetheless, despite its potential advantages in enhancing quantization accuracy, it significantly increases the number of scaling factors, posing another challenge for their optimizations via LSQ [32], which is widely adopted to optimize scaling factors in layer-wise quantized activations.
To address this limitation, we propose to adopt channel-wise migration on top of the layer-wise quantization for DW’s inputs. Specifically, as shown in Fig. 5, weights in DWConvs feature the same channel number as inputs, and each weight channel functions as an independent filter dedicated to processing its corresponding input channel. This arrangement enables us to assign distinct scaling factors to each weight channel. Thus, filter-wise quantization is essentially the channel-wise quantization for DW’s weights. Consequently, inter-channel variations of activations can be seamlessly transferred to weights through a mathematically equivalent transformation employing channel-wise migration factors :
where , , and are the output, input, weight, and migration factor for channel, respectively. donate the quantization function, and are the layer-wise and channel-wise scaling factors for and , respectively. This approach greatly facilitates activation quantization without hindering weight quantization. Note that weights can be pre-transformed before deployment to eliminate the on-chip computation. As for activations that depend on the input images during inference and thus cannot be pre-processed, we can fuse with to obtain a fused scaling factor in advance, thus avoiding the on-the-fly transformation:
Furthermore, the computation of can be formulated as:
where is the mean of the maximal values across channels. By comparing Figs. 3 (c)/(d) and Figs. 6 (a)/(b), it is evident that this approach can accomplish two essential objectives. Firstly, it compresses outliers, effectively reducing the value range of activations. Secondly, it amplifies smaller values, making them more amenable to quantization.
IV-C Filter-Wise Shifting for PW2’s Inputs
To eliminate the inter-channel asymmetries in PW2’s inputs, inspired by [33], we propose filter-wise shifting to pre-process PW2’s inputs before quantization. As expressed in Eq. (9), it subtracts each input channel by its channel-wise mean , thus obtaining the calibrated input centered around zero (see Fig. 6 (c)) and significantly reducing value ranges (see the comparison between Fig. 3 (f) and fig. 6 (d)). To accommodate the above filter-wise shifting for activations and keep the same functionality as the original PW, the original bias of the output channel needs to be updated to following Eq. (10), where donate the input channel number. This can be pre-computed to eliminate the on-chip processing.
IV-D Log2 Quantization for Divisors in MSAs
As illustrated in Figs. 4 (a) and (b), log2 quantization will allocate more bins to smaller values and vice versa. This inherent characteristic aligns with the algorithmic property of divisors in MSAs, where small values exhibit higher quantization sensitivity as explained in Sec. IV-A2. Thus, to improve the quantization resolution of small values in divisors of MSAs, we advocate adopting log2 quantization as follows:
where is log2-qauntized divisors, is the integer divisors generated by integer multiplications between and in Eq. (4), and and are their scaling factors. Note that can be pre-computed, while can be effectively implemented in the integer domain following [9]. Specifically, we first adopt the leading one detector (LOD) to find the index of the first non-zero bit of , then add with the value of the bit to obtain the result. For instance, if is , then the index of the first non-zero bit is and the value of the bit is , thus the log2-quantized () is .
By adopting log2 quantization for divisors, we can further replace hardware-unfriendly divisions in Eq. (4) with hardware-efficient bit-wise shifts, further enhancing hardware efficiency while boosting quantization performance.
V Trio-ViT’s Accelerator
V-A Design Considerations
To fully unleash our algorithmic benefits, developing a dedicated accelerator for quantized EfficientViT [12] is highly desired. However, this poses several challenges due to (i) various operation types within EfficientViT and (ii) the distinct computational pattern of its lightweight attention compared to the vanilla self-attention in standard ViTs [4, 5].
V-A1 Design Challenge # 1: Various Operation Types
As shown in Fig. 2 and introduced in Sec. III-A, there are mainly four types of operations in the Convolution-Transformer hybrid backbone of EfficientViT: generic convolutions (where output pixels are produced by the accumulation of partial sums within sliding windows along input channels), PWConvs (which essentially are generic convolutions with kernels), DWConvs (which process each input channel separately, thus only partial sums within the sliding window need to be accumulated), and matrix multiplications (MatMuls). For example, they account for , , , and of the total operation numbers in EfficientViT-B1, respectively, when input resolution is .
Design Choice # 1: Multipliers-Adder-Tree Architecture. Considering that PWConvs are the dominant type of operations, one natural thought is to build our dedicated accelerator with the Multipliers-Adder-Tree (MAT) architecture, which is a typical design to efficiently support PWConvs with channel parallelism [34, 35]. Specifically, as illustrated in Fig. 7 (a), each processing element (PE) lane in the MAT engine is responsible for the multiplication (via multipliers), summation (via the adder tree), and accumulation (via the accumulator) along the input channel dimension to generate each output pixel (for PWConvs) or partial sum (dubbed psum, for generic convolutions). Thus, the parallelism within the PE lane of the MAT engine is along the input channel dimension to facilitate psum reuse. Besides, inputs are broadcast to different PE lanes and multiplied with different weight filters to produce output/psum pixels from different output channels, thus the parallelism among PE lanes is along the output channel dimension to enhance input reuse.
Limitations. Although the MAT architecture can efficiently process PWConvs as well as easily support generic convolutions and MatMuls (which can be treated as PWConvs with large batch sizes), it has limited flexibility when handling DWConvs. Firstly, as for DWConvs, only psums generated from the same sliding window can be summed and accumulated, thus the achievable parallelism of each PE lane in MAT is limited by kernel sizes of DWConvs. Secondly, DWConvs in EfficientViT feature various kernel sizes ( and ) and strides ( and ), resulting in different sizes of sliding windows and distinct overlap patterns between adjacent sliding windows when conducting convolutions. This necessitates extra line buffers and substantial memory management overheads to support the multiplication and summation functionality within PE lanes for generating consecutive output pixels [34]. Thirdly, as there is no input reuse opportunity in DWConvs, this will lead to a high memory bandwidth requirement and otherwise low PE utilization when accommodating the output channel parallelism among PE lanes in the MAT architecture.
Design Choice # 2: Reconfigurable Architecture. To solve the above limitations, we can consider the reconfigurable architecture depicted in Fig. 7 (b), which incorporates multiple Reconfigurable Multiplier-ACcumulation units (R-MACs). (i) As depicted in Fig. 7 (c), when executing generic convolutions, PWConvs, and MatMuls, this architecture can be configured to operate in the down-forward accumulation mode to achieve the same functionality as the MAT architecture. This means that each PE lane supports the input channel parallelism to achieve psum reuse, while different PE lanes favor output channel parallelism to exploit the input reuse opportunity. (ii) As depicted in Fig. 7 (d), when executing DWConvs, R-MAC can be configured to run in the self-accumulation mode, thus partial sums within each sliding window can be temporally accumulated within each R-MAC, making this kind of design inherently supports DWConvs with various kernel sizes. Specifically, R-MACs within each PE lane can spatially compute multiple output pixels from different output channels to boost throughput, thus the parallelism is along the output channel here (see Fig. 7 (d) right).
Besides, weights can be broadcast to all PE lanes and multiplied with input pixels from different sliding windows to generate consecutive output pixels from the same output channel (see Fig. 7 (d) top left). By doing this, we can reuse overlaps among adjacent sliding windows with only several auxiliary registers and easily support DWConvs with different strides. For example, as shown in Fig. 7 (d) bottom left, where we take computations of the DWConv with a stride of executed on R-MACs arranged in the same row from PE lanes as the example. Initially, a sequence of input pixels is transmitted to input shift registers, which is then moved forward by cycles. During each cycle, the first pixels in shift registers are independently multiplied with the broadcast weight , generating the psums for consecutive output pixels. After cycles, the computation progresses to the next row of the input feature map and the corresponding filter, maintaining the same computation mode. This process is repeated until all rows are processed, resulting in output pixels. As for computations of DWConvs with a stride of , overlaps among adjacent sliding windows are spaced instead of successive. Thus, odd-column-indexed input pixels within each row are initially transmitted to shift registers for processing, followed by the even-column-indexed pixels. Weights also need to be broadcast following the same “first odd, then even” rule to accommodate this modified computation scheme. Thereby, the parallelism among PE lanes in this architecture is along the output feature map to enhance weight reuse.
Limitations. Despite its flexibility in supporting all types of operations in EfficientViT, there exist reconfigurable overheads. Firstly, the overheads in computational resources and buffers: each R-MAC needs a high-bit adder and a psum register to support the self-accumulation. Secondly, the overheads in control logic: extra multiplexers are required to simultaneously support both two accumulation patterns, i.e., self-accumulation and down-forward accumulation.
Our Proposed Design: Hybrid Architecture. Considering the fact that: (i) PWConvs are the dominant operations in EfficientViT [12] and the MAT architecture can efficiently support them, as well as (ii) EfficientViT incorporates various operation types, especially DWConvs with various kernel sizes and strides, and the R-MAC design can flexibly support them, we propose a hybrid architecture for our dedicated accelerator to marry the best of both worlds. Specifically, it consists of a MAT engine to efficiently process generic convolutions, PWConvs, and MatMuls, and a R-MAC engine to effectively support the above three operation types and DWConvs, thus enhancing flexibility while maintaining hardware efficiency.
Offered Opportunity: Inter-Layer Pipeline. Besides the efficiency and flexibility of our proposed hybrid architecture, it also offers an opportunity for the inter-layer pipeline, saving data access costs. Specifically, DWConvs are exclusively executed on the R-MAC engine in our hybrid accelerator and are sandwiched by two PWConvs in MBConvs of EfficientViT. Thus, when the R-MAC engine handles DWConvs, the resulting outputs can be immediately transmitted to the idle MAT engine and serve as inputs to participate in the computation of the subsequent PWConvs. This integration enables the computation of DWConvs and their following PWConvs to be fused, leading to enhanced hardware utilization and a reduction in data access costs from off-chip.
V-A2 Design Challenge # 2: Distinct Computational Pattern of Attention
As expressed in Eq. (4), after query and key undergo ReLU, there are five remaining steps for producing the final attention map : (i) MatMuls between and value ; (ii) toke-wise summation of ; then (iii) MatMuls of and outputs from step i; (iv) matrix-vector multiplications of and outputs from step ii; and finally (v) divisions between outputs from step iii (divisors) and step iv (dividends). Thanks to our log2 quantization for divisors as introduced in Sec. IV-D, the costly divisions can be substituted by hardware-efficient bit-wise shifts. Thus, it is evident that, besides multiplications, there are also element-wise summations and bit-wise shifts involved in computations of lightweight attention in EfficientViT. These multiplication-free element-wise operations are inherently incompatible with our multiplication-based PE arrays. Furthermore, they also exhibit low computational intensity, yielding increased bandwidth requirements or potential delays [36].
Our Proposed Solution: Low-Cost Auxiliary Processors. In addition to the MAT engine and R-MAC engine, which can be leveraged to effectively process MatMuls in the above steps i, iii, and iv, we further integrate several low-cost auxiliary processors into our hybrid architecture to facilitate the multiplication-free computations involved in lightweight attention. Particularly, we incorporate an adder tree to support the row-wise summation in step ii and a shifter array to handle the bit-wise shifts in step v. This architectural adjustment offers an opportunity for computation fusion within the attention (i.e., intra-layer pipeline, which will be explained in Sec. V-C), thus enhancing data reuse and easing bandwidth requirements.
Offered Opportunity: Intra-Layer Pipeline. Considering the MAT engine, R-MAC engine, and low-cost auxiliary processors in our hybrid design, the above computation steps in the attention that involve various operation types can be simultaneously handled by distinct computing units, thus offering the fusion opportunity. For example, (i) when in step i are executed on the MAT/R-MAT engine, can be broadcast to the auxiliary adder tree for performing the row-wise summation in step ii. Besides, (ii) when steps iii and iv are processed, their outputs can be immediately sent to the shifter array for element-wise divisions.
V-B Micro Architecture
As shown in Fig. 8, our dedicated accelerator consists of computing cores and several global buffers (buffer B/C and output buffer). Each computing core includes several internal buffers (buffer A and auxiliary/divisor buffers) and multiple computing units (R-MAC engine, MAT engine, auxiliary processors, and re-/log2 quantization modules). Specifically, as for computing units in each computing core, the R-MAC engine comprises PE lanes, each containing R-MACs and can be reconfigured to operate in either self-accumulation or down-forward accumulation modes. This flexibility enables the effective processing of all multiplication-based operations in EfficientViT, including generic convolutions, DWConvs, PWConvs, and MatMuls, as introduced in Design Choice # 2 in Sec. V-A1. The MAT engine consists of PE lanes, each including multipliers, and is developed to efficiently handle multiple multiplication-based operations in EfficientViT, excluding DWConvs, as explained in Design Choice # 1 in Sec. V-A1. Besides, the log2 quantization module is developed to quantize divisors in Eq. (4) following steps outlined at the end of first paragraph in Sec. IV-D, thus boosting quantization accuracy as well as enabling the conversion of costly divisions into hardware-efficient bit-wise shifts, as discussed in Sec. IV-D. Our accelerator is also equipped with several low-cost auxiliary processors, such as the adder tree and shifter array, to accommodate the computation of multiplication-free operations (e.g., row-wise summation and bit-wise shift) in MSAs, as discussed in Our Proposed Solution in Sec. V-A2. Additionally, a re-quantization module is also needed to re-quantize outputs following Eq. (12), where and are quantized output , input, and weight, respectively, and are their corresponding scaling factors, and / are both positive integers. By doing this, the floating-point re-scaling factors are converted into dyadic numbers (DN), and the re-quantization process can be implemented by integer-only multiplications and bit-wise shifts [11, 8, 7], thus facilitating both intra- and inter-layer pipelines on-chip.
As for internal buffers, buffer A broadcasts data to all PE lanes in the R-MAC engine and can also send data to the auxiliary adder tree. The auxiliary buffer cashes outputs from the R-MAC engine and broadcasts data to all PE lanes in the MAT engine, serving as a bridge between the two engines. The divisor buffer stores divisors in Eq. (4) and then transfers them to the log2 quantization module, in preparation for the following bit-wise shifts. Regarding global buffers, buffers B/C send data to all computing cores, where data are distributed and transmitted to different PE lanes in R-MAC/MAT engines. The output buffer stores data from all computing cores and directs them to the off-chip DRAM.
V-C Inter- and Intra-Layer Pipelines
As introduced above, our dedicated accelerator incorporates both multiplication-based engines (R-MAC and MAT engines, with DWConvs limited to the former) and multiplication-free engines (auxiliary processors that are designed to expedite computations in MSAs). This architecture inherently offers opportunities for pipeline processing, where various operations can be simultaneously executed on distinct computing units, thereby enhancing hardware utilization and throughput. As for the inter-layer pipeline, as shown in Figs. 9 (a) and (b), when the R-MAC engine handles DWConv, the resulting outputs are first subtracted by the channel-wise mean obtained on the calibration data to implement the filter-wise shifting introduced in Sec. IV-C, aiming to facilitate activation quantization. After that, they undergo re-quantization through the re-quantization module before being stored in the auxiliary buffer. Then, they are promptly directed to the idle MAT engine to serve as inputs for subsequent PWConv computations. Given that DWConvs entail fewer computations compared to PWConvs, once the processing of the current DWConv is completed, the R-MAC engine can be reassigned to participate in the concurrent computation of PWConv, alongside the MAT engine.
Regarding intra-layer pipeline, (i) when the R-MAC engine processes for the head, are broadcast to the auxiliary adder tree to generate the vector via row-wise summations. This implies that steps i/ii in the Sec. V-A2 can be executed simultaneously on distinct computing units. Concurrently, (ii) the MAT engine sequentially preform multiplications between and 1) already obtained as well as 2) to generate divisors and dividends in Eq. (4) for the head, respectively. This means that steps iv/iii in the Sec. V-A2 are computed consecutively on the MAT engine. During this process, the firstly generated divisors are cached in the divisor buffer and then routed to the log2 quantization module for log2 quantization. (iii) Once the dividends are obtained, they can be re-quantized via the re-quantization module and then sent to the auxiliary shifter array along with the already log2-quantized divisors to conduct element-wise divisions via bit-wise shifts, thus obtaining the final outputs of MSA. Note that once the computation for of all heads is finished, the R-MAC engine can be reused to compute divisors and dividends, together with the MAT engine.
VI Experimental Results
| Resources | BRAM | DSP | LUT | FF |
| Available | 912 | 2520 | 274080 | 548160 |
| Used | 162 (17.8%) | 1024 (40.6%) | 130628 (47.7%) | 152819 (27.9%) |
VI-A Experimental Setup
Dataset, Baselines, and Metrics. We validate our Trio-ViT’s post-training quantization algorithm on the ImageNet dataset [37]. Specifically, we randomly sample images from the training set as calibration data and then test on the validation set. ❶ To verify the effectiveness of our quantization engine, we consider seven baselines: MinMax, EMA [38], Percentile [39], OMSE [40], Bit-Split [41], EasyQuant [42], and the SOTA method FQ-ViT [9], for standard ViTs[4]/DeiTs[5], and compare with them in terms of accuracy (top-1 accuracy by default). ❷ To validate our dedicated accelerator, we consider nine baselines: (i) full-precision ViTs executed on the widely-used Edge GPU (NVIDIA Tegra X2); -bit ViTs quantized following (ii) FasterTransformer [43], (iii) I-BERT [44], and (iv) I-ViT [11], and accelerated on the Turing Tensor Core of GPU (NVIDIA 2080Ti), which supports efficient integer arithmetic via TVM; and full-precision EfficientViT executed on the (v) widely-used Edge CPU (Qualcomm Snapdragon 8Gen1 CPU) and Edge GPUs, including (vi) NVIDIA Jetson Nano and (vii) NVIDIA Jetson Orin; as well as two SOTA ViT accelerators, including (viii) the dedicated accelerator Auto-ViT-Acc [16] for standard ViTs and (ix) ViA [45] tailored for Swin Transformer [22], one of efficient ViTs. We compare them in terms of throughput, energy efficiency, frame rate, and DSP efficiency.
Accelerator Setup. Characteristics: The parallelism of computing engines in our accelerator (as depicted in Fig. 8) is configured to as . Thus, there are a total of multipliers in our accelerator, each can execute an -bit multiplication. To improve DSP utilization, we adopt the SOTA DSP packing strategy [46] to accommodate two -bit multiplications within each DSP, similar to Auto-ViT-Acc [16] for fair comparisons. Evaluation: We implement our accelerator with Verilog, synthesize through Vivado Design Suite, and evaluate on Xilinx ZCU102 FPGA at 200-MHz frequency. Table III lists our resource consumptions. Furthermore, we develop a cycle-accurate simulator for our accelerator to obtain fast and reliable estimations and verify them against the RTL implementation to ensure correctness.
| Method |
LinAttn |
| Param. (M) | – | – | 5.7 | 22.1 | 86.6 | 86.6 | ||||||||||
| GFLOPs | – | – | 1.3 | 4.6 | 17.6 | 17.6 | ||||||||||
| Full Precision | ✗ | ✗ | 72.21 | 79.85 | 81.85 | 84.53 | ||||||||||
| Base PTQ | ✗ | ✗ | 71.78 | 79.35 | 81.37 | 83.48 | ||||||||||
| Bit-Split [41] | - | 77.06 | 79.42 | - | ||||||||||||
| EasyQuant[42] | ✗ | ✗ | - | 76.59 | 79.36 | - | ||||||||||
| MinMax | 70.94 | 75.05 | 78.02 | 23.64 | ||||||||||||
| EMA [38] | 71.17 | 75.71 | 78.82 | 30.3 | ||||||||||||
| Percentile[39] | 71.47 | 76.57 | 78.37 | 46.69 | ||||||||||||
| OMSE [40] | 71.3 | 75.03 | 79.57 | 73.39 | ||||||||||||
| FQ-ViT [9] | ✓ | ✗ | 71.07 | 78.4 | 80.85 | 82.68 | ||||||||||
| Method | Softmax | LinAttn |
| Param. (M) | – | – | 9.1 | 9.1 | 9.1 | 24 | ||||||||||
| GFLOPs | – | – | 0.52 | 0.68 | 0.86 | 1.6 | ||||||||||
| Full Precision | ✗ | ✓ | 79.39 | 79.92 | 80.41 | 82.10 | ||||||||||
| Base PTQ | ✗ | ✓ | NaN | NaN | NaN | NaN | ||||||||||
| Ours | ✗ | ✓ | 78.64 | 78.93 | 79.58 | 80.97 | ||||||||||
R224 denotes the resolution of input images is , and so on.
VI-B Evaluation of Trio-ViT’s Post-Training Quantization
Results and Analysis. From Table IV, we can draw four conclusions. (i) The SOTA post-training quantization (PTQ) method FQ-ViT [9], which develops dedicated quantization schemes to fully quantize all operations in standard ViTs (including Softmax) to enhance hardware efficiency, suffers from accuracy compared to the full-precision models. Besides, it also yields an average accuracy compared to the base PTQ, where Softmax and other non-linear operations are not quantized, thus incurring non-negligible hardware costs. This demonstrates that the hardware-unfriendly non-linear operations are sensitive to quantization, hindering both the achievable hardware efficiency and quantization accuracy of standard ViTs. (ii) To address this limitation, the SOTA efficient ViT dubbed EfficientViT [12] has been proposed, which features Softmax-free linear attention (LinAttn) and can achieve much higher accuracy with even fewer parameters and computational costs. This underscores EfficientViT’s superiority, highlighting the need for quantization to facilitate its real-world applications. (iii) However, due to the distinct distributions of activation in MBConvs and MSAs, as introduced in Sec. IV-A, the vanilla PTQ method fails to quantize EfficientViT and even yields a Not-a-Number (NaN) issue. (iv) To solve this issue, we propose our dedicated PTQ engine, which can effectively quantize EfficientViTs with merely an average accuracy when compared with the full precision counterparts, demonstrating our effectiveness.
| MBConv Quant | MSA Quant | EfficientViT -B1-R224 | EfficientViT -B2-R224 | ||||
| Vanilla | CW Migration | FW Shifting | Uniform (8) | Log2 (4)* | |||
| – | – | – | – | – | 79.39 | 82.10 | |
| ✓ | ✓ | NaN | NaN | ||||
| ✓ | – | – | 3.23 | 0.68 | |||
| ✓ | – | – | 7.51 | 78.52 | |||
| ✓ | – | – | 28.75 | 0.94 | |||
| ✓ | ✓ | – | – | 79.05 | 81.36 | ||
| ✓ | ✓ | ✓ | NaN | NaN | |||
| ✓ | ✓ | ✓ | 78.64 | 80.97 | |||
denotes the 8-bit uniform quantization and 4-bit log2 quantization.
Effectiveness of Our Dedicated Quantization Engine. As shown in Table V, we can see that: As for the quantization within MBConvs, (i) due to the inter-channel variations in DW’s inputs and inter-channel asymmetries in PW2’s inputs, as introduced in Sec. IV-A1, vanilla uniform quantization fails to quantize MBConvs in EfficientViT. (ii) By incorporating our channel-wise migration and filter-wise shifting, which are proposed to solve the above two issues, respectively, and are both indispensable, we can effectively quantize MBConvs with only an average accuracy. On top of this, regarding the quantization of lightweight MSA, (iii) owing to the extreme quantization sensitivity of smaller values in divisors, as illustrated in Sec. IV-A2, the vanilla -bit uniform quantization yields a NaN issue. (iv) Thus, we advocate adopting log2 quantization for divisors, which assigns more bins to smaller values and is inherently compatible with the algorithmic property of divisors, thus allowing for the effective quantization of divisors with a mere -bit.
| CW Quantization | 69.25 | 78.83 | 79.36 | 79.98 | ||||||||||
| CW Migration | 78.64 | 78.93 | 79.58 | 80.97 | ||||||||||
| Improve (%) | 9.39 | 0.10 | 0.22 | 0.99 | ||||||||||
Effectiveness of Proposed Channel-Wise Migration. Considering the unique algorithmic property of DWConvs, where each channel of weights serves as an independent filter to process each input channel, channel-wise quantization serves as a straightforward solution to solve the inter-channel variations in DW’s input, as explained in Sec. IV-B. However, it will greatly increase the number of scaling factors, challenging quantization optimization via LSQ [32] and limiting achievable accuracy. Thus, we advocate adopting channel-wise migration on top of layer-wise quantization. As shown in Table VI, by doing this, we can offer an average accuracy, demonstrating our superiority in solving the inter-channel variations of DW’s input while maintaining optimization efficiency.
VI-C Evaluation of Trio-ViT’s Dedicated Accelerator
Comparisons with SOTA Baselines on GPUs/CPUs. We follow [47, 26] to scale up the hardware resources of our accelerator to have comparable peak throughput with the general computing platform (i.e., NVIDIA 2080Ti GPU) to enable fair comparisons with GPUs/CPUs. As shown in Fig. 10 (where the y-axis is plotted on a logarithmic scale for better illustration), we can achieve much better hardware efficiency compared to SOTA baselines on GPUs/CPUs, validating our effectiveness. Specifically, (i) when compared with the full-precision DeiTs [5] on Edge GPU (Tegra X2), we can achieve and throughput and energy efficiency, respectively. (ii) As for comparisons with 8-bit DeiTs quantized by FasterTransformer [43], I-BERT [44] and I-ViT [11], and executed on the Turning Tensor Core of the NVIDIA 2080Ti GPU, we can offer and throughput and energy efficiency, respectively. Furthermore, (iii) regarding the comparison with full precision EfficientViT [12] on edge CPU, we can gain up to and throughput and energy efficiency, respectively. (iv) When compared with EfficientViT on edge GPUs (NVIDIA Jetson Nano and Jetson Orin), we can gain and in terms of throughput and energy efficiency.
| Accelerator | Via [45] | Auto-ViT-Acc [16] | Ours | |||
| Device | Xilinx | Xilinx | Xilinx | |||
| Alveo U50 | ZCU102 | ZCU102 | ||||
| Frequency (MHz) | 300 | 150 | 200 | |||
| Format | FP16 | FIX8 | FIX8 | |||
| Model | Swin-T | Deit-Small | Deit-Base | EfficientViT | EfficientViT | EfficientViT |
| -R224 | -R224 | -R224 | -B2-R288 | -B1-R224 | -B2-R256 | |
| DSP Used | 2420 | 1936 | 2066 | 1024 | 1024 | 1024 |
| Latency (ms) | 14.5 | 12.8 | 38.6 | 1.12 | 2.02 | 2.64 |
| FPS | 68.8 | 78.1 | 25.9 | 895 | 494 | 379 |
| Throughput | 310 | 711 | 900 | 769 | 791 | 796 |
| DSP Efficiency | 0.13 | 0.37 | 0.44 | 0.75 | 0.77 | 0.78 |
| (GOPS/DSP) | ||||||
| Energy Efficiency | 7.92 | 84.1 | 95.7 | 105 | 108 | 109 |
| (GOPS/W) | ||||||
| Accuracy (%) | 81.3 | 79.69 | 81.93 | 79.58 | 80.97 | 81.62 |
Comparisons with SOTA ViT Accelerators. From Table VII, we can see that: (i) due to the superiority of our quantization engine and the promising hardware efficiency of EfficientViT compared to standard DeiT, we gain the lowest latency and highest frame rate (FPS) under comparable accuracy. Particularly, we can achieve up to FPS compared to Via [45], a dedicated accelerator for an efficient ViT dubbed Swin-Transformer-Tiny (Swin-T) [22], as well as FPS over the dedicated ViT/DeiT accelerator Auto-ViT-Acc [16]. (ii) Furthermore, owing to the hybrid design of our dedicated accelerator, which is proposed to effectively support various operators in the Convolution-Transformer hybrid architecture of EfficientViT with enhanced hardware utilization, we can achieve the highest hardware utilization efficiency. For example, we can offer up to and DSP efficiency when compared with Via and Auto-ViT-Acc, respectively. (iii) Additionally, thanks to our developed pipeline architecture aiming to facilitate both inter- and intra-layer fusion, we gain the best energy efficiency, i.e., up to and compared to Via and Auto-ViT-Acc, respectively.
VII Conclusion
In this paper, we have proposed, developed, and validated Trio-ViT, the first post-training quantization and acceleration framework dedicated to the state-of-the-art (SOTA) efficient Vision Transformer (ViT), dubbed EfficientViT. Specifically, at the algorithm level, we propose a tailored post-training quantization engine that incorporates several innovative quantization schemes to effectively quantize EfficientViT with enhanced quantization accuracy. At the hardware level, we develop a dedicated accelerator integrating a hybrid design and a pipeline architecture, aiming to boost hardware efficiency. Extensive experimental results consistently prove our effectiveness. Particularly, we can gain up to and frame rate with comparable accuracy over state-of-the-art ViT accelerators.
Limitations and Future Work. It has been widely demonstrated that model layers exhibit varying degrees of sensitivity to quantization, thus allocating the same bit to all layers is deemed sub-optimal in both accuracy and efficiency [20, 18]. Therefore, our future research will focus on exploring mixed quantization, considering variations in both quantization bits and schemes (such as fix-point and power-of-two).